Roberts Company has the following information available for the current year:

Standard:

Material 4.25 feet per unit @ $2.75 per foot

Labor 7 direct labor hours @9.25 per unit

Actual:

Material 128000 feet used (130000 feet were purchased costing $364000)

Labor 212000 direct labor hours incurred costing $1971600

30000 units were produced

A. Refer to Roberts Company. Compute the labor rate variance and labor efficiency variances. Clearly indicate if the variances are favorable or unfavorable.

B. Refer to Roberts Company. Compute the material price variance and material quantity variances. Clearly indicate if the variances are favorable or unfavorable.
Answer #1

A

Labor rate variances :

(Actual hour × Actual rate) - (Actual hour × standard rate)

Or actual labor cost - Standard cost for actual hour

Actual labor cost for 30,000 unit=$1,971,600

Standard labor cost for 30,000 unit

= Standard rate× actual hour used

=9.25×212,000=$1,961,000

Labor rate variances = 1,971,600 - 1,961,000

= $10600(unfavorable)

Labor efficiency variances

= (Actual hour ×st.rate) - (St. Hour × st. Rate)

Total standard hour for 30,000unit

7×30,000 = 210,000

Labor efficiency variances =

(212000×9.25) - (210,000×9.25)

=$18500(unfavorable)

B

Material price variances :

( Actual quantity ×Actual rate) -(Actual quantity × st.rate)

Or

Actual quantity at actual rate - Actual quantity at st. Rate

Material used 128,000feet but purchases 130,000 feet

So cost of 128,000feet= 364,000×128000/130,000

=$358,400

Actual quantity at std rate = 128,000×2.75=$352,000

Material price variances =

358,400- 352,000 = $6400 ( Unfavourable)

Material Quantity variances =

(Actual. Quantity × Std rate)-(std. Quantity× std rate)

(128000×2.75) - (30,000×4.25×2.75)

352,000 -350,625= 1375 Unfavourable

..